Zeronknight (zeronknight):

Write 38% as a ratio in lowest terms. * 19 to 100 38 to 50 50 to 19 19 to 50

Zeronknight (zeronknight):

@sammixboo

sammixboo (sammixboo):

It wouldn't be 19 to 100, because that would be 19%. We can put 38% into the form of 38/100, and then find the greatest common factor for 38/100

sammixboo (sammixboo):

Simplify 38/100, in other terms.

Zeronknight (zeronknight):

Uh how?

Zeronknight (zeronknight):

XD. I really have no idea on 38/100.

OpenStudy (abdullahm):

% means over 100 38% = 38/100 You can then start by dividing the numerator by the greatest common factor. To find the greatest common factor, list all the factors of the number. For example, the factors of 100 are: 100: 1, 2, 4, 5, 10, 20, 25, 50, 100 Do the same for 38, and tell me what you get :)

Zeronknight (zeronknight):

1, 38? lol

sammixboo (sammixboo):

That's fine! Let's write out all the numbers you can multiply to get 38 and 100. 38 = 1, 2, 19, 38 100 =1, 2, 4, 5, 10, 20, 25, 50, 100

Zeronknight (zeronknight):

Oh 2

OpenStudy (abdullahm):

So, divide the numerator and denominator both by 2 and you'll get your final answer!

sammixboo (sammixboo):

Right, now divide 38 by 2, and 100 by 2. \(\dfrac{38 \div 2}{100\div 2}\)

sammixboo (sammixboo):

Yes

Zeronknight (zeronknight):

38 / 2 = 19 100 / 2 = 50

Zeronknight (zeronknight):

So D?

OpenStudy (abdullahm):

Yes

OpenStudy (lpruitt4):

Bill will be a good scientist. will be isunderlined a. action verb b. begin verb